### v1

@@ Abstract
Newton’s method may exhibit slower convergence than vanilla Gradient Descent in its initial phase on strongly convex problems. Classical Newton-type multilevel methods mitigate this but, like Gradient Descent, achieve only linear convergence near the minimizer. We introduce an adaptive multilevel Newton-type method with a principled automatic switch to full Newton once its quadratic phase is reached. The local quadratic convergence for strongly convex functions with Lipschitz continuous Hessians and for self-concordant functions is established and confirmed empirically. Although per-iteration cost can exceed that of classical multilevel schemes, the method is efficient and consistently outperforms Newton’s method, Gradient Descent, and the multilevel Newton method, indicating that second-order methods can outperform first-order methods even when Newton’s method is initially slow.
@@ 1 Introduction
Among optimization methods applied for solving (1), the pure Newton method has the most powerful theory, exhibiting quadratic convergence rate in a neighborhood of the minimizer. Once Newton’s method enters this neighborhood, convergence is extremely fast, doubling the number of the correct digits at each iteration. In practice this means that highly accurate solutions can be reached in only a few iterations, i.e., three or four are typically enough [3, 27]. Pure Newton’s main drawback is related to the cost of forming the Newton direction and the storage requirements of the Hessian matrix. The per-iteration cost of Newton’s method scales cubically with the problem dimension $n$, which can significantly affect the overall runtime of the algorithm. Nonetheless, this cost is often justified by the method’s rapid convergence once the iterates are sufficiently close to the solution $x^{*}$. In particular, the local quadratic convergence of Newton’s method is especially advantageous in solving highly ill-conditioned problems where high-precision solutions are required—tasks that may be infeasible for methods limited to linear convergence rates.
In contrast, the global convergence of (damped) Newton’s method is less efficient. Specifically, for strongly convex functions, the damped Newton method achieves only a linear convergence rate [21]. This rate matches that of Gradient Descent for the same class of functions, rendering the higher computational cost of forming and solving systems involving the Hessian matrix unjustified when compared to the substantially cheaper iterations of Gradient Descent. The issue of whether second-order methods are indeed advantageous at the initial stages of optimization was also raised in [25] for the purposes of the accelerated Cubic Newton method. Moreover, empirical observations further support this discrepancy, with Gradient Descent often exhibiting faster progress than second-order methods during the initial stages of the minimization process [2]. Consequently, due to its significantly lower per-iteration cost, Gradient Descent may substantially outperform Newton’s method in terms of runtime when the iterates are far from the minimizer $x^{*}$.
@@
To the best of our knowledge, a robust hybrid strategy that automatically and adaptively determines a meaningful transition point between methods in a principled manner is not available in the literature. In this paper we propose an Adaptive Multilevel (AML) Newton method to address this gap in the literature. For the proposed method, we establish a local quadratic convergence rate for both strongly convex functions with Lipschitz continuous Hessians and self-concordant functions, respectively. To achieve these rates, we introduce a new criterion whereby the proposed method adaptively alternates between $m>1$ subspaces (or coarse levels) of $\mathbb{R}^{n}$, selecting at each iteration the smallest one that provides a theoretically justified and efficient decrease in the norm of the gradients. Specifically, the theory suggests that the AML–Newton method automatically selects very small coarse levels (i.e., faster iterates) in early stages where the rate of convergence is typically slow, and larger ones near $x^{*}$ to obtain its quadratic rate. The local quadratic rates are easy to interpret and are comparable to that of Newton’s method. As in Newton’s method, once the method enters its quadratic phase, it will maintain this rate until the end of the minimization process. Furthermore, we establish a connection with random subspace Newton-type methods. In particular, for the Random Subspace Newton (RSN) method, we show that it achieves a local quadratic rate with high probability, based on random embeddings that satisfy the Johnson–Lindenstrauss (JL) lemma [19].
We recall that local quadratic convergence rates are generally not applicable to classical multilevel (ML–Newton) or random subspace Newton-type methods. Without additional structural assumptions on problem (1), one can typically expect only a local linear-quadratic rate [11, 16, 17, 33]. In this work, we significantly improve the local convergence behavior of these methods through a simple modification: we alternate between subspaces by looking one step ahead in the minimization process, monitoring gradient progress within the relevant subspaces. This strategy ensures local quadratic convergence as the gradient is provably shown to decay more rapidly within the subspaces.
@@
The multilevel framework, originally developed for solving linear elliptic partial differential equations (see [1] for a review), was brought into optimization to accelerate the convergence of second-order methods such as trust-region and Newton methods. MG/OPT [24] was the first to extend the PDE multigrid framework to optimization. Building on this, [13] introduces a fully recursive trust-region scheme (RMTR) across discretization levels with global guarantees, and [14] strengthens the theory with global convergence under milder assumptions on coarse models and transfer operators while allowing inexact subproblem solves. For constrained problems, [41] develops an adaptive multilevel inexact SQP framework. In the unconstrained setting, [17] provides convergence guarantees for a multilevel Newton method in the strongly convex case, while [33] extends these results to self-concordant functions with bounds that do not depend on unknown problem constants; [36] further analyzes a low-rank variant with theory for self-concordant objectives. Moreover, [35] introduces a multilevel framework for regularized Newton methods in nonconvex optimization. This work is closest to ours: it shows that the multilevel trial steps can match the convergence guarantees of regularized Newton, but its practical efficiency has yet to be verified empirically.
Beyond their well-established theoretical foundations, multilevel methods also deliver efficient performance in modern, real-world applications. Multilevel training for deep learning includes globally convergent RMTR-style methods for ResNets [22], stochastic MG/OPT for ResNets with reported speedups [37], multilevel minimization tailored to depth hierarchies [10], and a stochastic preconditioned gradient method for escaping saddle points and flat areas efficiently [32]. Furthermore [12] reframes multilevel as block-coordinate descent, and demonstrates gains for deep-learning PDE surrogates, while [4] proposes multilevel Levenberg–Marquardt training scheme that optimizes neural networks. Layer-parallel (MGRIT) approaches expose parallelism across layers for ResNets [15] and GRUs [23], while multigrid-style training schedules accelerate video models [40]. Recent software demonstrates scalable layer-parallel training on MPI/GPU systems [8], and multilevel-in-width schemes regularize and speed network training [28].
@@
An example of these variants is Randomized Block Coordinate Descent (RBCD) where at each iteration it updates a randomly selected subset of coordinates and therefore requires only the corresponding gradient components. In Newton–type block variants, one additionally uses the associated block of the Hessian, see [26, 30, 39]. Under standard assumptions (smoothness, strong convexity), Block Coordinate Descent enjoys global convergence with linear rate; the constants are usually better for the Newton variant than for first-order RBCD because the step uses the curvature in the active block.
Random subspace methods generalize block coordinate methods by replacing coordinate-aligned blocks with arbitrary low-dimensional subspaces specified by a sketch matrix $R$. In practice, $R$ often originates from the Johnson–Lindenstrauss (JL) lemma [19], which effectively reduces the problem’s dimensionality without discarding essential second-order information. The RSN [11] efficiently uses sketching to compute a Newton step restricted to a low-dimensional subspace. SDNA [29] is a stochastic dual Newton method for empirical risk minimization that updates small blocks of dual variables using local Hessian information. The Randomized Block Cubic Newton method [9] performs cubic regularized Newton updates on randomly sampled variable blocks using block Hessians to capture curvature while keeping iterations cheap. A general randomized subspace framework to tackle large non-convex problems yielding cheaper iterations with global first-order convergence guarantees is proposed in [5]. The work in [6] applies the ideas of the general subspace framework [5] to Adaptive Regularization with Cubics (ARC) to achieve scalable cubic regularization steps in low-dimensional random subspaces, while preserving ARC’s optimal first- and second-order complexity guarantees.
@@
    This paper significantly advances the theory of multilevel methods such as RMTR [13] and ML–Newton [17, 33]. Our key difference is a new acceptance rule that predicts when a coarse step will be effective, providing a principled mechanism to switch automatically from a subspace method to full Newton steps precisely when they are most efficient.
@@ 2 Notation and Preliminaries
We say that a symmetric matrix $A\in\mathbb{R}^{n\times n}$ is positive semi-definite if $x^{\top}Ax\geq 0$ for all $x\in\mathbb{R}^{n}$ and we write $A\succeq 0$. Similarly, it is positive definite if $x^{\top}Ax>0$ for all $x\in\mathbb{R}^{n}\setminus\{0\}$ and we write $A\succ 0$. We write $A\succeq B$ (resp. $A\succ B$) if ABA-B is positive semi-definite (resp. positive definite). Moreover, $I_{d}$ denotes the $d\times d$ identity matrix. For a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the gradient and Hessian at $x\in\mathbb{R}^{n}$ are denoted by $\nabla f(x)$ and $\nabla^{2}f(x)$, respectively. In addition, the standard Euclidean norm is denoted by $\|x\|$. For a matrix $P\in\mathbb{R}^{m\times n}$ with rank $n<m$, we denote by $\sigma_{1}(P)\geq\sigma_{2}(P)\geq\cdots\geq\sigma_{n}(P)$ its singular values as $\sigma_{i}(P)\coloneq\sqrt{\lambda_{i}(P^{\top}P)}$, where $\lambda_{i}(P^{\top}P)>0$ is the $i$-th largest eigenvalue of the positive definite matrix $P^{\top}P$. Further, we have $\|P\|=\|P^{\top}\|=\sigma_{1}(P)$, for more details on matrix analysis see [18].
@@ 2.1 Newton’s Method for Strongly Convex and Self-Concordant Functions
Let $\mu,L>0$ such that $f$ is $\mu$-strongly convex with $L$-Lipschitz continuous Hessians. Then, pure Newton’s steps achieve a local quadratic convergence [3] with rate given by
@@
If $M=1$ in (3), the function $f$ is called standard self-concordant. This class is important because it guarantees convergence rates that are independent of unknown problem constants. For more details on the convergence rate of Newton’s method under strong convexity and self-concordance we refer the reader to the books [3, 27].
@@ 2.2 The Classical Multilevel Framework for Newton’s Method
The general multilevel approach for unconstrained optimization has been explored in various settings, see for instance [13, 24, 38]. Subsequent works studied it as a two-level method, arising naturally from the structure of Newton’s method when constructing its iterates [17, 33]. In this section, we unify these perspectives and present a general framework that incorporates a finite hierarchy of subspaces, tailored specifically to Newton-type methods. We refer to this approach as the Multilevel Newton (ML–Newton) method.
@@
i.e., the model $\psi_{i-1}$ is first-order coherent w.r.t. the model $f_{i}$, see also [24]. To ensure its second-order coherency, we define $f_{i-1}$ as
@@
The choice of $f_{i-1}$ in (7) is standard in multilevel methods, and is referred to as a Galerkin model, originally introduced in [38]. The work in [38] and subsequent ones, e.g., [17, 33], have extensively evaluated its performance, demonstrating its effectiveness in practical applications. Within the multilevel framework, the model $\psi_{i-1}$ is referred to as coarse, since $n_{i}<n$, whereas $f_{m}$ (or equivalently $f$) is termed the fine model.
@@
To conclude this section, it remains to discuss a procedure for selecting an appropriate level, or coarse model, based on which the point $x_{k+1}$ will be generated. Ideally, one would select $\psi_{1}$ to obtain $d_{1,k}$, since $n_{1}\leq n_{i}$ for all $i\geq 2$ and thus, $d_{1,k}$ is more efficient to compute. However, coarse directions may not always be effective. This may occur when $\tilde{R}_{i}\nabla f_{i}(x_{i,k})=0$ while $\nabla f_{i}(x_{i,k})\neq 0$. This is possible since it might be that $\nabla f_{i}(x_{i,k})\in\operatorname{null}\big({\tilde{R}_{i}}\big)$. Classical multilevel methods evaluate the effectiveness of a coarse model $\psi_{i-1}$ before computing the associated coarse direction. This assessment can be carried out efficiently by checking the following condition (see also [13]): let $\gamma\in\big(0,\min\big\{1,\min_{i}\big\|\tilde{R}_{i}\big\|\big\}\big)$ and $\varepsilon\in(0,1)$ — then the coarse model $\psi_{i-1}$ is selected if

### v2

@@ Abstract
Newton’s method may exhibit slower convergence than vanilla Gradient Descent in its initial phase on strongly convex problems. Classical Newton-type multilevel methods mitigate this but, like Gradient Descent, achieve only linear convergence near the minimizer. We introduce an adaptive multilevel Newton-type method with a principled automatic switch to full Newton once its quadratic phase is reached. The local quadratic convergence for strongly convex functions with Lipschitz continuous Hessians and for self-concordant functions is established and confirmed empirically. Although per-iteration cost can exceed that of classical multilevel schemes, the method is efficient and consistently outperforms Newton’s method, Gradient Descent, and the multilevel Newton method, indicating that second-order methods can outperform first-order methods even when Newton’s method is initially slow. The promising empirical results open new avenues for designing reduced-cost second- and high-order methods with extremely fast convergence rates.
@@ 1 Introduction
Among optimization methods applied for solving (1), the pure Newton method has the most powerful theory, exhibiting quadratic convergence rate in a neighborhood of the minimizer. Once Newton’s method enters this neighborhood, convergence is extremely fast, doubling the number of the correct digits at each iteration. In practice this means that highly accurate solutions can be reached in only a few iterations, i.e., three or four are typically enough [boyd2004convex, nesterov2018lectures]. Pure Newton’s main drawback is related to the cost of forming the Newton direction and the storage requirements of the Hessian matrix. The per-iteration cost of Newton’s method scales cubically with the problem dimension $n$, which can significantly affect the overall runtime of the algorithm. Nonetheless, this cost is often justified by the method’s rapid convergence once the iterates are sufficiently close to the solution $x^{*}$. In particular, the local quadratic convergence of Newton’s method is especially advantageous in solving highly ill-conditioned problems where high-precision solutions are required—tasks that may be infeasible for methods limited to linear convergence rates.
In contrast, the global convergence of (damped) Newton’s method is less efficient. Specifically, for strongly convex functions, the damped Newton method achieves only a linear convergence rate [karimireddy2018global]. This rate matches that of Gradient Descent for the same class of functions, rendering the higher computational cost of forming and solving systems involving the Hessian matrix unjustified when compared to the substantially cheaper iterations of Gradient Descent. The issue of whether second-order methods are indeed advantageous at the initial stages of optimization was also raised in [nesterov2008accelerating] for the purposes of the accelerated Cubic Newton method. Moreover, empirical observations further support this discrepancy, with Gradient Descent often exhibiting faster progress than second-order methods during the initial stages of the minimization process [bottou2018optimization]. Consequently, due to its significantly lower per-iteration cost, Gradient Descent may substantially outperform Newton’s method in terms of runtime when the iterates are far from the minimizer $x^{*}$.
@@
To the best of our knowledge, a robust hybrid strategy that automatically and adaptively determines a meaningful transition point between methods in a principled manner is not available in the literature. In this paper we propose an Adaptive Multilevel (AML) Newton method to address this gap in the literature. For the proposed method, we establish a local quadratic convergence rate for both strongly convex functions with Lipschitz continuous Hessians and self-concordant functions, respectively. To achieve these rates, we introduce a new criterion whereby the proposed method adaptively alternates between $m>1$ subspaces (or coarse levels) of $\mathbb{R}^{n}$, selecting at each iteration the smallest one that provides a theoretically justified and efficient decrease in the norm of the gradients. Specifically, the theory suggests that the AML–Newton method automatically selects very small coarse levels (i.e., faster iterates) in early stages where the rate of convergence is typically slow, and larger ones near $x^{*}$ to obtain its quadratic rate. The local quadratic rates are easy to interpret and are comparable to that of Newton’s method. As in Newton’s method, once the method enters its quadratic phase, it will maintain this rate until the end of the minimization process. Furthermore, we establish a connection with random subspace Newton-type methods. In particular, for the Random Subspace Newton (RSN) method, we show that it achieves a local quadratic rate with high probability, based on random embeddings that satisfy the Johnson–Lindenstrauss (JL) lemma [johnson1984extensions].
We recall that local quadratic convergence rates are generally not applicable to classical multilevel (ML–Newton) or random subspace Newton-type methods. Without additional structural assumptions on problem (1), one can typically expect only a local linear-quadratic rate [gower2019rsn, hanzely2020stochastic, ho2019newton, tsipinakis2024multilevel]. In this work, we significantly improve the local convergence behavior of these methods through a simple modification: we alternate between subspaces by looking one step ahead in the minimization process, monitoring gradient progress within the relevant subspaces. This strategy ensures local quadratic convergence as the gradient is provably shown to decay more rapidly within the subspaces.
@@
The multilevel framework, originally developed for solving linear elliptic partial differential equations (see [borzi2009multigrid] for a review), was brought into optimization to accelerate the convergence of second-order methods such as trust-region and Newton methods. MG/OPT [nash2000multigrid] was the first to extend the PDE multigrid framework to optimization. Building on this, [gratton2008recursive] introduces a fully recursive trust-region scheme (RMTR) across discretization levels with global guarantees, and [gross2009convergence] strengthens the theory with global convergence under milder assumptions on coarse models and transfer operators while allowing inexact subproblem solves. For constrained problems, [ziems2011adaptive] develops an adaptive multilevel inexact SQP framework. In the unconstrained setting, [ho2019newton] provides convergence guarantees for a multilevel Newton method in the strongly convex case, while [tsipinakis2024multilevel] extends these results to self-concordant functions with bounds that do not depend on unknown problem constants; [tsipinakis2023lowrank] further analyzes a low-rank variant with theory for self-concordant objectives. Moreover, [tsipinakis2025regularized] introduces a multilevel framework for regularized Newton methods in nonconvex optimization. This work is closest to ours: it shows that the multilevel trial steps can match the convergence guarantees of regularized Newton, but its practical efficiency has yet to be verified empirically.
Beyond their well-established theoretical foundations, multilevel methods also deliver efficient performance in modern, real-world applications. Multilevel training for deep learning includes globally convergent RMTR-style methods for ResNets [kopanicakova2022globally], stochastic MG/OPT for ResNets with reported speedups [von2021training], multilevel minimization tailored to depth hierarchies [gaedke2021multilevel], and a stochastic preconditioned gradient method for escaping saddle points and flat areas efficiently [tsipinakis2023simba]. Furthermore [gratton2024block] reframes multilevel as block-coordinate descent, and demonstrates gains for deep-learning PDE surrogates, while [calandra2019approximation] proposes multilevel Levenberg–Marquardt training scheme that optimizes neural networks. Layer-parallel (MGRIT) approaches expose parallelism across layers for ResNets [gunther2020layer] and GRUs [moon2022parallel], while multigrid-style training schedules accelerate video models [wu2020multigrid]. Recent software demonstrates scalable layer-parallel training on MPI/GPU systems [cyr2025torchbraid], and multilevel-in-width schemes regularize and speed network training [ponce2023multilevel].
@@
An example of these variants is Randomized Block Coordinate Descent (RBCD) where at each iteration it updates a randomly selected subset of coordinates and therefore requires only the corresponding gradient components. In Newton–type block variants, one additionally uses the associated block of the Hessian, see [nesterov2012efficiency, richtarik2014iteration, wright2015coordinate]. Under standard assumptions (smoothness, strong convexity), Block Coordinate Descent enjoys global convergence with linear rate; the constants are usually better for the Newton variant than for first-order RBCD because the step uses the curvature in the active block.
Random subspace methods generalize block coordinate methods by replacing coordinate-aligned blocks with arbitrary low-dimensional subspaces specified by a sketch matrix $R$. In practice, $R$ often originates from the Johnson–Lindenstrauss (JL) lemma [johnson1984extensions], which effectively reduces the problem’s dimensionality without discarding essential second-order information. The RSN [gower2019rsn] efficiently uses sketching to compute a Newton step restricted to a low-dimensional subspace. SDNA [qu2016sdna] is a stochastic dual Newton method for empirical risk minimization that updates small blocks of dual variables using local Hessian information. The Randomized Block Cubic Newton method [doikov2018randomized] performs cubic regularized Newton updates on randomly sampled variable blocks using block Hessians to capture curvature while keeping iterations cheap. A general randomized subspace framework to tackle large non-convex problems yielding cheaper iterations with global first-order convergence guarantees is proposed in [cartis2022randomised]. The work in [cartis2025random] applies the ideas of the general subspace framework [cartis2022randomised] to Adaptive Regularization with Cubics (ARC) to achieve scalable cubic regularization steps in low-dimensional random subspaces, while preserving ARC’s optimal first- and second-order complexity guarantees.
@@
    This paper significantly advances the theory of multilevel methods such as RMTR [gratton2008recursive] and ML–Newton [ho2019newton, tsipinakis2024multilevel]. Our key difference is a new acceptance rule that predicts when a coarse step will be effective, providing a principled mechanism to switch automatically from a subspace method to full Newton steps precisely when they are most efficient.
@@ 2 Notation and Preliminaries
We say that a symmetric matrix $A\in\mathbb{R}^{n\times n}$ is positive semi-definite if $x^{\top}Ax\geq 0$ for all $x\in\mathbb{R}^{n}$ and we write $A\succeq 0$. Similarly, it is positive definite if $x^{\top}Ax>0$ for all $x\in\mathbb{R}^{n}\setminus\{0\}$ and we write $A\succ 0$. We write $A\succeq B$ (resp. $A\succ B$) if ABA-B is positive semi-definite (resp. positive definite). Moreover, $I_{d}$ denotes the $d\times d$ identity matrix. For a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the gradient and Hessian at $x\in\mathbb{R}^{n}$ are denoted by $\nabla f(x)$ and $\nabla^{2}f(x)$, respectively. In addition, the standard Euclidean norm is denoted by $\|x\|$. For a matrix $P\in\mathbb{R}^{m\times n}$ with rank $n<m$, we denote by $\sigma_{1}(P)\geq\sigma_{2}(P)\geq\cdots\geq\sigma_{n}(P)$ its singular values as $\sigma_{i}(P)\coloneq\sqrt{\lambda_{i}(P^{\top}P)}$, where $\lambda_{i}(P^{\top}P)>0$ is the $i$-th largest eigenvalue of the positive definite matrix $P^{\top}P$. Further, we have $\|P\|=\|P^{\top}\|=\sigma_{1}(P)$, for more details on matrix analysis see [horn2012matrix].
@@ 2.1 Newton’s Method for Strongly Convex and Self-Concordant Functions
Let $\mu,L>0$ such that $f$ is $\mu$-strongly convex with $L$-Lipschitz continuous Hessians. Then, pure Newton’s steps achieve a local quadratic convergence [boyd2004convex] with rate given by
@@
If $M=1$ in (3), the function $f$ is called standard self-concordant. This class is important because it guarantees convergence rates that are independent of unknown problem constants. For more details on the convergence rate of Newton’s method under strong convexity and self-concordance we refer the reader to the books [boyd2004convex, nesterov2018lectures].
@@ 2.2 The Classical Multilevel Framework for Newton’s Method
The general multilevel approach for unconstrained optimization has been explored in various settings, see for instance [gratton2008recursive, nash2000multigrid, MR2587737]. Subsequent works studied it as a two-level method, arising naturally from the structure of Newton’s method when constructing its iterates [ho2019newton, tsipinakis2024multilevel]. In this section, we unify these perspectives and present a general framework that incorporates a finite hierarchy of subspaces, tailored specifically to Newton-type methods. We refer to this approach as the Multilevel Newton (ML–Newton) method.
@@
i.e., the model $\psi_{i-1}$ is first-order coherent w.r.t. the model $f_{i}$, see also [nash2000multigrid]. To ensure its second-order coherency, we define $f_{i-1}$ as
@@
The choice of $f_{i-1}$ in (7) is standard in multilevel methods, and is referred to as a Galerkin model, originally introduced in [MR2587737]. The work in [MR2587737] and subsequent ones, e.g., [ho2019newton, tsipinakis2024multilevel], have extensively evaluated its performance, demonstrating its effectiveness in practical applications. Within the multilevel framework, the model $\psi_{i-1}$ is referred to as coarse, since $n_{i}<n$, whereas $f_{m}$ (or equivalently $f$) is termed the fine model.
@@
To conclude this section, it remains to discuss a procedure for selecting an appropriate level, or coarse model, based on which the point $x_{k+1}$ will be generated. Ideally, one would select $\psi_{1}$ to obtain $d_{1,k}$, since $n_{1}\leq n_{i}$ for all $i\geq 2$ and thus, $d_{1,k}$ is more efficient to compute. However, coarse directions may not always be effective. This may occur when $\tilde{R}_{i}\nabla f_{i}(x_{i,k})=0$ while $\nabla f_{i}(x_{i,k})\neq 0$. This is possible since it might be that $\nabla f_{i}(x_{i,k})\in\operatorname{null}\big({\tilde{R}_{i}}\big)$. Classical multilevel methods evaluate the effectiveness of a coarse model $\psi_{i-1}$ before computing the associated coarse direction. This assessment can be carried out efficiently by checking the following condition (see also [gratton2008recursive]): let $\gamma\in\big(0,\min\big\{1,\min_{i}\big\|\tilde{R}_{i}\big\|\big\}\big)$ and $\varepsilon\in(0,1)$ — then the coarse model $\psi_{i-1}$ is selected if